# From Autoencoders to CycleGAN: Robust Unpaired Face Manipulation via Adversarial Learning

### v1

@@ I Introduction
In recent years, deepfake applications have brought face-changing technologies into the mainstream, demonstrating the ability of generative models to produce highly realistic human faces. Facial manipulation tasks generally fall into three main categories: face synthesis, face swapping, and facial attribute/expression editing. From a technical standpoint, these applications are built upon advances in deep generative modeling, including autoencoders, generative adversarial networks (GANs), and adversarial architectures for high-quality image-to-image translation such as pix2pix (paired) and CycleGAN (unpaired) [5, 19].
@@
This project addresses these limitations by moving from autoencoder-based facial manipulation to an improved CycleGAN framework that leverages adversarial learning for more robust and flexible results. Concretely, we incorporate spectral normalization to constrain the discriminator’s Lipschitz behavior and stabilize training [13]; identity- and perceptual-guided learning (frozen face encoder and VGG-Face-style features) to preserve subject identity and high-level structure [11]; semantic/landmark-weighted cycle consistency to maintain facial geometry; and lightweight architectural refinements—patchwise contrastive regularization for local detail, multi-scale discriminators, and optional attention in the generator—to reduce artifacts and improve coherence [7]. We further adopt practical training refinements (TTUR, EMA, and differentiable augmentations) that improve convergence smoothness and sample efficiency.
@@ II Related Work
Deep generative modeling has progressed along three principal lines: autoregressive models (AR), variational autoencoders (VAEs), and generative adversarial networks (GANs). Autoregressive approaches factorize the joint density into a product of conditional distributions, which yields tractable likelihoods but can be slow at sample time due to their sequential nature [14]. VAEs introduce a probabilistic latent space with amortized inference [9, 10]. Given $x$, the encoder produces parameters $(\mu(x),\sigma(x))$ of a Gaussian approximate posterior $q_{\phi}(z\!\mid\!x)$, and reparameterization draws
@@
Extensions improve perceptual fidelity by comparing in learned feature spaces [11] and by blending variational inference with adversarial training [1].
GANs formulate generation as a two-player minimax game between a generator $G$ and a discriminator $D$ [4]. The generator maps latent noise $z\sim p_{z}$ (e.g., standard Gaussian) to the data space, producing $G(z)$, while the discriminator receives either real data $x\sim p_{\text{data}}$ or synthetic samples $G(z)$ and outputs the probability of being real. The standard value function is
@@
Training alternates gradient updates to $D$ and $G$; at equilibrium, the model distribution $p_{g}$ matches $p_{\text{data}}$ and $D$ cannot discriminate better than chance. A large body of subsequent work explores improved architectures and training objectives, including conditional GANs for paired translation [5], feature-metric regularization for autoencoding [11], hybrid variational–adversarial models such as CVAE-GAN [1], and spectral normalization for stabilizing training [13]. State-of-the-art unconditional generators (e.g., StyleGAN) have further advanced image quality via style-based modulation and progressive growing [7].
CycleGAN extends adversarial learning to unpaired image–to–image translation by learning two mappings between domains $A$ and BB—a forward generator $G\!:\!A\!\to\!B$ and a backward generator $F\!:\!B\!\to\!A$—together with discriminators $D_{A}$ and $D_{B}$ that judge whether translated images look like real samples from their respective domains [19]. During training, the discriminators encourage each generator to produce outputs that are indistinguishable from target-domain images (adversarial learning), while a cycle-consistency constraint requires that translating to the other domain and back approximately recovers the original input. This combination preserves content and layout without paired supervision, which is especially useful in face manipulation where pose, expression, and illumination differ across domains. For practical use, widely adopted PyTorch implementations are available [18, 17].
Recent diffusion models generate high-fidelity images by denoising in a learned latent space, dramatically improving sample quality and controllability [15]. Although diffusion is beyond the scope of our main method, we include it for completeness as a contemporary alternative to GAN-based approaches.
@@ III Dataset and Features
We extracted $\sim$1,000 face frames from the 2024 U.S. Presidential Debate video using FFmpeg [16] for model training and validation ($\sim$500 per identity). We also utilize VGGFace2 [2] for large-scale pose/diversity coverage and reference the CelebA attributes dataset [12] as a standard benchmark for identity/attribute supervision. In addition, we experimented with domain-biased content using the Anime Face Dataset from Kaggle [6] to assess cross-domain transferability. These datasets allow evaluating both within-domain latent operations (VAE) and cross-domain translation (CycleGAN). Figure 2 shows an example of face manipulation via deepfake-style techniques [3].
@@ IV-A Problem Setup and Baseline
We consider unpaired translation between two facial domains $X$ and $Y$ with data distributions $p_{X}$ and $p_{Y}$. Following CycleGAN [19], we learn generators $G_{X\to Y}$ and $G_{Y\to X}$ and discriminators $D_{Y}$ and $D_{X}$ trained adversarially [4] with cycle-consistency. For $X\!\to\!Y$ the adversarial objective is
@@
As in [19], we use ResNet-based generators (9 residual blocks), PatchGAN discriminators, and optimize with Adam [8]. Our generators are image-to-image mappings (they take an image from the other domain as input rather than white noise), which is standard for CycleGAN.
@@ IV-B Spectral Normalization and Lipschitz Control
We regularize the generator and discriminator via spectral normalization (SN) [13]. The discriminator’s function class critically affects GAN stability; constraining it to be (approximately) Lipschitz-continuous prevents gradient explosions and improves boundedness of statistics. Under the original GAN objective, the optimal discriminator has the form
@@ IV-C Proposed Improvements for Face Manipulation
Beyond pixels, we include a feature/perceptual loss [11] using a frozen backbone $\Phi$ (e.g., VGG-Face features):
@@
We insert lightweight self-attention at $1/16$ and $1/8$ resolutions in $G$ to capture long-range dependencies (hairlines, jaw contours) and adopt multi-scale discriminators $D^{1},D^{2}$ (full and half resolution) to detect artifacts at different scales. Inspired by style-based generators [7], we add channel-wise modulation (AdaIN) in late residual blocks to stabilize style transfer.
@@
Alongside SN [13], we use global-norm gradient clipping, an exponential moving average (EMA) of generator weights, and the Two-Time-Scale Update Rule (TTUR): two $D$ updates per $G$ update early in training. We apply differentiable data augmentation (translation, color jitter, cutout) to $D$’s inputs, use mixed-precision training (FP16), and progressively resize inputs ($128\!\to\!256$). For completeness, latent diffusion [15] provides strong generative priors, but we keep our method fully GAN-based for efficiency. Also, for hybrid supervisonion that when a small paired subset is available (e.g., pose-aligned frames), we add a conditional loss in the spirit of pix2pix [5] for those minibatches, while retaining the unpaired objectives elsewhere.
@@
We set $\lambda_{\mathrm{cyc}}{=}10$, $\lambda_{\mathrm{id}}{=}1$, $\lambda_{\mathrm{perc}}{=}0.5$, $\lambda_{\mathrm{sem}}{=}2$, $\lambda_{\mathrm{lmk}}{=}1$, and $\lambda_{\mathrm{con}}{=}0.5$ (tuned on a dev set). Optimization uses Adam [8] with $\beta_{1}{=}0.5$, $\beta_{2}{=}0.999$, initial LR $2{\times}10^{-4}$ with linear decay, batch size 2288 (GPU-dependent), EMA decay $0.999$, and TTUR (two $D$ steps per $G$ step for the first 25% of training, then 1:11{:}1).
@@ V Experimental Setup
We build upon public PyTorch implementations [18, 17]. The network architecture is shown in Figure 3. Generators use 9-block ResNets with skip connections at $1/4$ and $1/2$ scales, lightweight self-attention at $1/16$ and $1/8$, and AdaIN in late blocks [7]. Discriminators are PatchGANs with spectral normalization [13], evaluated at two scales (full and half resolution). Identity/perceptual features are computed with a frozen face encoder and VGG-Face-style layers (feature-space losses motivated by [11]). Face parsing masks and landmarks are precomputed and treated as constants. Differentiable augmentations are applied to $D$’s inputs; mixed precision (FP16) is enabled throughout.
@@
VAE: convolutional encoder/decoder, latent dimension $128$, Adam (lr $1{\times}10^{-4}$). CycleGAN: official PyTorch repo [18, 17]; 9-block ResNet generators, PatchGAN discriminators, Adam (lr $2{\times}10^{-4}$) with linear decay. pix2pix (paired baseline): U-Net generator, PatchGAN discriminator, Adam (lr $2{\times}10^{-4}$).

### v2

@@ I Introduction
In recent years, deepfake applications have brought face-changing technologies into the mainstream, demonstrating the ability of generative models to produce highly realistic human faces. Facial manipulation tasks generally fall into three main categories: face synthesis, face swapping, and facial attribute/expression editing. From a technical standpoint, these applications are built upon advances in deep generative modeling, including autoencoders, generative adversarial networks (GANs), and adversarial architectures for high-quality image-to-image translation such as pix2pix (paired) and CycleGAN (unpaired) [pix2pix, zhu2017cyclegan].
@@
This project addresses these limitations by moving from autoencoder-based facial manipulation to an improved CycleGAN framework that leverages adversarial learning for more robust and flexible results. Concretely, we incorporate spectral normalization to constrain the discriminator’s Lipschitz behavior and stabilize training [spectralnorm]; identity- and perceptual-guided learning (frozen face encoder and VGG-Face-style features) to preserve subject identity and high-level structure [larsen2016aebp]; semantic/landmark-weighted cycle consistency to maintain facial geometry; and lightweight architectural refinements—patchwise contrastive regularization for local detail, multi-scale discriminators, and optional attention in the generator—to reduce artifacts and improve coherence [karras2019stylegan]. We further adopt practical training refinements (TTUR, EMA, and differentiable augmentations) that improve convergence smoothness and sample efficiency.
@@ II Related Work
Deep generative modeling has progressed along three principal lines: autoregressive models (AR), variational autoencoders (VAEs), and generative adversarial networks (GANs). Autoregressive approaches factorize the joint density into a product of conditional distributions, which yields tractable likelihoods but can be slow at sample time due to their sequential nature [ostrovski2018autoregressive]. VAEs introduce a probabilistic latent space with amortized inference [Kingma2014, kingma2014vae]. Given $x$, the encoder produces parameters $(\mu(x),\sigma(x))$ of a Gaussian approximate posterior $q_{\phi}(z\!\mid\!x)$, and reparameterization draws
@@
Extensions improve perceptual fidelity by comparing in learned feature spaces [larsen2016aebp] and by blending variational inference with adversarial training [bao2017cvaegan].
GANs formulate generation as a two-player minimax game between a generator $G$ and a discriminator $D$ [gan]. The generator maps latent noise $z\sim p_{z}$ (e.g., standard Gaussian) to the data space, producing $G(z)$, while the discriminator receives either real data $x\sim p_{\text{data}}$ or synthetic samples $G(z)$ and outputs the probability of being real. The standard value function is
@@
Training alternates gradient updates to $D$ and $G$; at equilibrium, the model distribution $p_{g}$ matches $p_{\text{data}}$ and $D$ cannot discriminate better than chance. A large body of subsequent work explores improved architectures and training objectives, including conditional GANs for paired translation [pix2pix], feature-metric regularization for autoencoding [larsen2016aebp], hybrid variational–adversarial models such as CVAE-GAN [bao2017cvaegan], and spectral normalization for stabilizing training [spectralnorm]. State-of-the-art unconditional generators (e.g., StyleGAN) have further advanced image quality via style-based modulation and progressive growing [karras2019stylegan].
CycleGAN extends adversarial learning to unpaired image–to–image translation by learning two mappings between domains $A$ and BB—a forward generator $G\!:\!A\!\to\!B$ and a backward generator $F\!:\!B\!\to\!A$—together with discriminators $D_{A}$ and $D_{B}$ that judge whether translated images look like real samples from their respective domains [zhu2017cyclegan]. During training, the discriminators encourage each generator to produce outputs that are indistinguishable from target-domain images (adversarial learning), while a cycle-consistency constraint requires that translating to the other domain and back approximately recovers the original input. This combination preserves content and layout without paired supervision, which is especially useful in face manipulation where pose, expression, and illumination differ across domains. For practical use, widely adopted PyTorch implementations are available [cyclegan_repo, cyclegan_github].
Recent diffusion models generate high-fidelity images by denoising in a learned latent space, dramatically improving sample quality and controllability [rombach2022ldm]. Although diffusion is beyond the scope of our main method, we include it for completeness as a contemporary alternative to GAN-based approaches.
@@ III Dataset and Features
We extracted $\sim$1,000 face frames from the 2024 U.S. Presidential Debate video using FFmpeg [ffmpeg] for model training and validation ($\sim$500 per identity). We also utilize VGGFace2 [vggface2] for large-scale pose/diversity coverage and reference the CelebA attributes dataset [liu2015celeba] as a standard benchmark for identity/attribute supervision. In addition, we experimented with domain-biased content using the Anime Face Dataset from Kaggle [animeface_kaggle] to assess cross-domain transferability. These datasets allow evaluating both within-domain latent operations (VAE) and cross-domain translation (CycleGAN). Figure 2 shows an example of face manipulation via deepfake-style techniques [deepfake].
@@ IV-A Problem Setup and Baseline
We consider unpaired translation between two facial domains $X$ and $Y$ with data distributions $p_{X}$ and $p_{Y}$. Following CycleGAN [zhu2017cyclegan], we learn generators $G_{X\to Y}$ and $G_{Y\to X}$ and discriminators $D_{Y}$ and $D_{X}$ trained adversarially [gan] with cycle-consistency. For $X\!\to\!Y$ the adversarial objective is
@@
As in [zhu2017cyclegan], we use ResNet-based generators (9 residual blocks), PatchGAN discriminators, and optimize with Adam [adam]. Our generators are image-to-image mappings (they take an image from the other domain as input rather than white noise), which is standard for CycleGAN.
@@ IV-B Spectral Normalization and Lipschitz Control
We regularize the generator and discriminator via spectral normalization (SN) [spectralnorm]. The discriminator’s function class critically affects GAN stability; constraining it to be (approximately) Lipschitz-continuous prevents gradient explosions and improves boundedness of statistics. Under the original GAN objective, the optimal discriminator has the form
@@ IV-C Proposed Improvements for Face Manipulation
Beyond pixels, we include a feature/perceptual loss [larsen2016aebp] using a frozen backbone $\Phi$ (e.g., VGG-Face features):
@@
We insert lightweight self-attention at $1/16$ and $1/8$ resolutions in $G$ to capture long-range dependencies (hairlines, jaw contours) and adopt multi-scale discriminators $D^{1},D^{2}$ (full and half resolution) to detect artifacts at different scales. Inspired by style-based generators [karras2019stylegan], we add channel-wise modulation (AdaIN) in late residual blocks to stabilize style transfer.
@@
Alongside SN [spectralnorm], we use global-norm gradient clipping, an exponential moving average (EMA) of generator weights, and the Two-Time-Scale Update Rule (TTUR): two $D$ updates per $G$ update early in training. We apply differentiable data augmentation (translation, color jitter, cutout) to $D$’s inputs, use mixed-precision training (FP16), and progressively resize inputs ($128\!\to\!256$). For completeness, latent diffusion [rombach2022ldm] provides strong generative priors, but we keep our method fully GAN-based for efficiency. Also, for hybrid supervisonion that when a small paired subset is available (e.g., pose-aligned frames), we add a conditional loss in the spirit of pix2pix [pix2pix] for those minibatches, while retaining the unpaired objectives elsewhere.
@@
We set $\lambda_{\mathrm{cyc}}{=}10$, $\lambda_{\mathrm{id}}{=}1$, $\lambda_{\mathrm{perc}}{=}0.5$, $\lambda_{\mathrm{sem}}{=}2$, $\lambda_{\mathrm{lmk}}{=}1$, and $\lambda_{\mathrm{con}}{=}0.5$ (tuned on a dev set). Optimization uses Adam [adam] with $\beta_{1}{=}0.5$, $\beta_{2}{=}0.999$, initial LR $2{\times}10^{-4}$ with linear decay, batch size 2288 (GPU-dependent), EMA decay $0.999$, and TTUR (two $D$ steps per $G$ step for the first 25% of training, then 1:11{:}1).
@@ V Experimental Setup
We build upon public PyTorch implementations [cyclegan_repo, cyclegan_github]. The network architecture is shown in Figure 3. Generators use 9-block ResNets with skip connections at $1/4$ and $1/2$ scales, lightweight self-attention at $1/16$ and $1/8$, and AdaIN in late blocks [karras2019stylegan]. Discriminators are PatchGANs with spectral normalization [spectralnorm], evaluated at two scales (full and half resolution). Identity/perceptual features are computed with a frozen face encoder and VGG-Face-style layers (feature-space losses motivated by [larsen2016aebp]). Face parsing masks and landmarks are precomputed and treated as constants. Differentiable augmentations are applied to $D$’s inputs; mixed precision (FP16) is enabled throughout.
@@
VAE: convolutional encoder/decoder, latent dimension $128$, Adam (lr $1{\times}10^{-4}$). CycleGAN: official PyTorch repo [cyclegan_repo, cyclegan_github]; 9-block ResNet generators, PatchGAN discriminators, Adam (lr $2{\times}10^{-4}$) with linear decay. pix2pix (paired baseline): U-Net generator, PatchGAN discriminator, Adam (lr $2{\times}10^{-4}$).